OpenStudy (anonymous):

"Gamma radiation has a frequency of 2.50 * 1022 l/s. What is the energy, in joules, of the photon?"

OpenStudy (aaronq):

\[E=h \nu \]

OpenStudy (anonymous):

Formula E=hv Units E = J h = J * s v = 1/s Example: E = ? (J) h = 6.626E-34 (J*s) v = 2.50E22 (1/s) \[6.626\times 10^{ -34 }(J.s)*2.50\times 10^{ 22 }(1/s)=1.656\times 10^{ 49 }(J)\] Energy of one mole of photons:1.656*10^ 49 Joules/mol photon 1 mol photon: 6.022E23 molecules photon \[\frac { 1.656\times 10^{ 49 }J }{ 1\quad mol\quad photon } *\frac { 1\quad mol\quad photon }{ 6.022x10^{ 23 }\quad molecule\quad photon } =2.749x10^{ 25 } Joules/molecule\ photon\quad \] Energy of a photon: 2.749E25 J/molecule photon